OpenStudy (anonymous):

Simplify the expression. (csc^2xsec^2x)/(sec^2x+csc^2x)

OpenStudy (anonymous):

replace cosec x by 1/(sin x) and sec x by 1/(cos x) and then simplify u may get 1 as the final answer

OpenStudy (anonymous):

Okay I did and got, (1/sin^2)(1/cos^2x) all over 1/cos^2x + 1/sin^2x

OpenStudy (anonymous):

I get confused from there

OpenStudy (anonymous):

@matricked What do I do from here?

OpenStudy (anonymous):

= (1/sin^2)(1/cos^2x) / ( 1/cos^2x + 1/sin^2x ) = (1/sin^2)(1/cos^2x) / ( ( sin^2x + cos^2x )/ (cos^2x *sin^2x) ) = (1/sin^2)(1/cos^2x) * (cos^2x *sin^2x) / ( ( sin^2x + cos^2x ) ) =1 we know sin^2x + cos^2x = 1

OpenStudy (anonymous):

On the second row, how did you get rid of the 1 over the denominators in 1/cos^2x+1/sin^2x?

OpenStudy (anonymous):

Any help please?

OpenStudy (anonymous):

1/cos^2x+1/sin^2 = ( sin^2x + cos^2x )/ (cos^2x sin^2x) ) [taking LCM and adding]

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

It just looks like you made the denominator into the numerator and vice versa

OpenStudy (anonymous):

can you calculate (1/2+ 1/3)

OpenStudy (anonymous):

Yes, 5/6

OpenStudy (anonymous):

did you ever gt in btwn like (3+2)/(3*2)

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

does it looks like you made the denominator into the numerator and vice versa

OpenStudy (anonymous):

similar steps are done for 1/cos^2x+1/sin^2x = ( sin^2x + cos^2x )/ (cos^2x sin^2x) )

OpenStudy (anonymous):

No

OpenStudy (anonymous):

OHHH, thank you

OpenStudy (anonymous):

ok what abt solving 1/a + 1/b

OpenStudy (anonymous):

I got it now, I finally understand now, I feel really dumb

OpenStudy (anonymous):

you should ask if you don't understand